You are here: Foswiki>Commissioning Web>Com20120403 (03 Apr 2012, AndrewRakich)Edit Attach

UT 20120403 Laser Tracker Collimation

Observer: DThompson (Tucson), ARakich (LBTO)
Telescope Operator: DGH (LBTO)
Telescope Support: (LBTO)
Instrument Support: JMorris (LBTO)

Summary

Got a good set of data after a shakey start.

Details

Telescope turned over to Andrew and I at 11:00 UT Seeing 2".

Found bug in Dave's RB code that was subtracting laser Tracker (LT) vector after range bal completed. The nice thing is this in itself gave a good confirmation that the LT was performing well, as subsequent active optics drove in exactly the missing range bal vector, +/- a few 10's of microns.

9:37 RB bug proved intracible, will troubleshoot during the day, moving on to laser tracker accuracy tests.

9:45 First, taking globals of M2 and putting them on M1 so I can easily enter my M2 change vectors as M2 globals.

9:46 Collimating on BS9155.

9:52 Measuring Laser Tracker "nominal points" on collimated mirror positions.

9:55 entered LT vector 025300 , collimated, got -0.022 -0.201 -0.005 20.953 -2.248

10:00 sending acquire preset to BS9158 EL 82 degrees.

10:02 running LT script and getting new vector for M2 global offsets.

10.07 IE 15 CA 4.1 (from IE -34 CA -2.9)

10:07, forgot to clear active optics before starting, so new "collimated" active optics values should subtract ones noted at 9:55.

10:12 OK, realised a few things about this "new plan" Have to clear global offsets and active optrics before each new preset. Have to manually correct astigmatism first before sending preset.

10:19 entered vector 31457 and now collimating to see what active optics residuals are... 0.869 0.283 -0.047 and corresponding coma free Rx Ry

10:24 Acquire preset to BS9150 El 37.

10:26 correcting pointing and astigmatism manually

10:27 laser tracker measurement. entering vector 32658 into M2 global offsets.

10:35 active optics converged to -0.295 -0.162 -0.130 16.8 -30.7

10:36 el 74 vector 33830 active optics converged to 0.463 0.187 -0.276 -19.478 48.258 FWHM after coma and focus removed 2.5" (before, 1.6")

10:57 EL 38.8 vector 35846 ActO -> -0.426 -0.111 -.496 11 -44

11:08 BS9158 EL 55 RESET NOMINAL POINTS as suspect mirror has only just reached equilibrium

11:19 after resetting nominals to collimated telescope, deteleted all offsets and remeasured with LT to see if we recover alignment. After laser tracker vector 41908, ActO-> 0.020 -0.065 -0.053 6.7 2.1

11:27 BS9166 EL 84 LT vector 42627, ActO-> 0.547 0.459 0.041 -48 57

11:34 BS9148 EL 31 LT vector 43707, ActO-> -0.578 -0.082 0.067 8.5 -60

11:46 BS9162 EL 69 LT vector 44736, ActO-> 0.644 0.397 -0.087 -41.3 67.1

11:48 BS9153 EL 40 LT vector 45804, ActO-> -0.432 -0.075 -0.046 7.8 -45

12:06 BS9170 EL 56 LT vector 50906, ActO-> -.477 -.252 -.191 -0.006 0.015 -0.001968

12:44 Got lots of data, looking into it...

-- DavidThompson - 03 Apr 2012
Topic revision: r5 - 03 Apr 2012, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback